### v1

# Real-is-Sim: Bridging the Sim-to-Real Gap with a Dynamic Digital Twin for Real-World Robot Policy Evaluation

Abstract
Recent advancements in behavior cloning have enabled robots to perform complex manipulation tasks. However, accurately assessing training performance remains challenging, particularly for real-world applications, as behavior cloning losses often correlate poorly with actual task success. Consequently, researchers resort to success rate metrics derived from costly and time-consuming real-world evaluations, making the identification of optimal policies and detection of overfitting or underfitting impractical. To address these issues, we propose real-is-sim, a novel behavior cloning framework that incorporates a dynamic digital twin (based on Embodied Gaussians) throughout the entire policy development pipeline: data collection, training, and deployment. By continuously aligning the simulated world with the physical world, demonstrations can be collected in the real world with states extracted from the simulator. The simulator enables flexible state representations by rendering image inputs from any viewpoint or extracting low-level state information from objects embodied within the scene. During training, policies can be directly evaluated within the simulator in an offline and highly parallelizable manner. Finally, during deployment, policies are run within the simulator where the real robot directly tracks the simulated robot’s joints, effectively decoupling policy execution from real hardware and mitigating traditional domain-transfer challenges. We validate real-is-sim on the PushT manipulation task, demonstrating strong correlation between success rates obtained in the simulator and real-world evaluations. Videos of our system can be found at https://realissim.rai-inst.com.
I INTRODUCTION
Figure 1: The real-is-sim framework, illustrating the information flow between its components. A policy trained in a correctable physics simulator (leveraging Embodied Gaussians) controls a simulated robot. Real-world observations continuously update the simulator, maintaining its state close to ground truth. The real robot then mirrors the simulated robot’s joint positions. This approach shifts the sim-to-real gap challenge from the policy to the adaptable physics simulator.
Simulation environments offer a highly favorable setting for developing and evaluating robotic manipulation policies. They support continuous monitoring of task performance during training, enable large-scale parallel rollouts, and allow for precise control over resets and environment variation. Crucially, simulation also affords flexible and often privileged access to state information, such as exact object poses or custom visual observations from arbitrary viewpoints, that can simplify policy learning. These advantages make behavior cloning particularly effective in simulation. However, transferring these benefits to the real world remains challenging. Real-world policy evaluation is slow, labor-intensive, and limited in scale, often relying on training loss as a proxy for success despite its poor correlation with actual task performance [1, 2, 3]. This gap between simulation and real-world applicability motivates the core question driving our work: How can we retain the strengths of simulation for behavior cloning while operating in the real world?
Despite the clear advantages of simulation, integrating these benefits into real-world policy development remains difficult. Broadly, two paradigms exist. The first, real-only training, relies exclusively on physical hardware [4]. While this offers immediate applicability, since policies are trained and deployed in the same environment, it comes with significant limitations: (i) it prevents offline evaluation, which is essential for diagnosing failure modes and tracking performance, and (ii) it restricts large-scale parallelization, which is useful for exploration, hyperparameter tuning, and proactive planning. The second paradigm, sim-to-real, aims to overcome these constraints by training in simulation and transferring policies to the real world. This allows for extensive offline evaluation and experimentation at scale [5]. However, it introduces the well-known sim-to-real gap: discrepancies in observations and dynamics between simulation and reality that can degrade policy performance. Bridging this gap often requires extensive, task-specific tuning of either the simulator or the policy, efforts that are difficult to generalize or scale.
Current approaches treat simulation and reality as separate domains, creating a mismatch between the observational inputs a learned manipulation policy receives during training and deployment. This discrepancy also necessitates ensuring that the simulated robot behaves identically to its real-world counterpart, making policy transfer from simulation to reality challenging. We propose real-is-sim, a paradigm shift where a correctable simulator remains always-in-the-loop, even during real-world deployment. A correctable simulator synchronizes with the physical world using sensor data (in our case, RGB images), creating a live, dynamic digital twin of the robot’s environment. The first simulator to achieve this generally and in real-time is Embodied Gaussians [6]. Embodied Gaussians combines the use of Gaussian splatting and a particle-based physics system to represent both the visual and physical aspects of the world. It allows, through its differentiable rendering process, the correction of the simulator state from observations received from cameras. In the real-is-sim framework, the always-in-the-loop simulator acts as a mediator between the physical world and the policy. Policies are trained exclusively on representations derived from the simulator’s state, effectively transforming real-world interactions into simulated states. This approach collapses the traditional divide between training and deployment domains.
Real-is-sim collects synchronized simulation states during real-world interactions, which serve as demonstration trajectories for policy training. Policies are trained on arbitrary representations extracted from these simulated states (which may include object poses, robot configuration, and/or rendered images from fixed or robot-mounted viewpoints). The policies output desired joint-space actions that control the simulated robot. Critically, this process is identical in both online and offline operational modes. In offline mode, the simulator operates without real-world coupling, executing the policy in a purely virtual environment. In online mode, the real robot acts as a follower and tracks the simulated robot’s joint positions, while real-world sensor data (RGB cameras) continuously correct the simulator’s state. This design (illustrated in Figure 1) ensures the policy remains fully decoupled from the physical hardware — it interacts solely with the simulated robot and environment in all modes. As such it allows the policies to always operate on in-distribution simulator states, even when deployed in the real world.
By maintaining a persistent, real-time digital twin of the physical environment, real-is-sim unifies simulation and reality into a single, coherent interface for policy development. This operational symmetry across both offline and online modes unlocks several key benefits: consistent policy behavior between simulation and deployment, scalable and safe offline evaluation, and flexible, perception-agnostic representations for policy learning. In this work, we show how real-is-sim fundamentally rethinks the role of simulation, not just as a training tool, but as a medium for end-to-end policy evaluation and deployment. We validate our framework on the PushT manipulation task, demonstrating that offline simulation evaluations closely predict real-world performance. Real-is-sim offers a scalable path forward for efficient training, reliable assessment, and robust deployment of robotic policies in the real world.
Figure 2: Comparison of three paradigms: Real-is-Sim, Real-Only, and Sim-to-Real, highlighting their online (real-world interaction) and offline (simulation-based) capabilities. Real-is-Sim offers a unified framework where the offline mode is a simplified version of the online, lacking only real-time correction. This ensures seamless transferability. Real-Only is confined to real-world execution. Sim-to-Real struggles with distribution mismatch and dynamics discrepancies, making successful transfer uncertain. In the Real-is-Sim paradigm, successful transfer between online and offline modes is mostly dependent on whether environment synchronization was successful during the demo collection phase.
II RELATED WORK
II-A Real World Policy Evaluation
Evaluating robotic policies directly in the real world remains the most reliable method for assessing performance in practical tasks [4]; however, it is often labor-intensive, time-consuming, and requires substantial human oversight [1]. In many experimental pipelines, researchers must manually reset the environment, initiate policy execution, and supervise task completion. For example, in the benchmark introduced by Cheng et al. [7], the authors report manually aligning a T-shaped object to match one of 20 predefined start poses for each policy rollout. This reset process involved visual overlays to ensure pose accuracy and was repeated across all baselines. This process is not parallelizable and poses a scalability bottleneck when evaluating across multiple tasks, policies, and environment variations.
To mitigate these challenges, a growing body of work has turned to evaluating real-world policies in simulation environments, even when those policies are trained on real data [8, 9, 10, 11]. This strategy offers safety, repeatability, and especially scalability, enabling parallelized policy assessment across diverse conditions. Prior works have spent large amounts of effort trying to improve the correlation between simulated and real world performance by improving visual appearance [8, 11], refining object models [8, 10], modeling control latency [9], and tracking cross-corelation between object classes [11].
In contrast to these evaluation-centric approaches, our framework incorporates the simulator as an integral component throughout the entire policy development pipeline: training, evaluation, and deployment. Rather than constructing simulation for post hoc policy assessment, we instead continuously synchronize the simulation with real-world sensor data during execution. This enables policies trained entirely in simulation to be deployed directly on real robots without additional fine-tuning or adaptation. By maintaining real-time alignment between the simulated and physical environments, our method ensures that the policy operates within a unified and consistent observation and control space. This paradigm shift facilitates seamless transfer between simulation and reality, enabling scalable evaluation and robust deployment under a single, coherent framework.
Figure 3: The two modes of the Real-is-Sim framework: online and offline. In online mode, the real robot operates alongside the correctable simulator (Embodied Gaussians). Synchronization is maintained through fictitious visual forces derived from the rendering loss between observed and real images, as well as by the real robot mimicking the simulated robot’s state. In offline mode, parallel environments execute the learned policy and evaluate its success rate. The simulator can run up to 7x faster than real time with policy sampling in the loop.
II-B Bridging the Sim-to-Real Gap for Robotic Manipulation
As discussed above, apart from direct real world evaluation [4], simulation of the corresponding real world environment offers a promising alternative. A major challenge lies in the discrepancy between the simulated states and reality particularly. To bridge this gap, numerous strategies have been proposed. Domain randomization injects variability into the simulation (e.g., textures, lighting, physics) to force robustness. Early work by Tobin et al. [12] demonstrated zero-shot sim-to-real transfer using randomized visuals. While powerful, this approach often requires excessive variability and long training times. Domain adaptation methods attempt to align visual or latent representations between sim and real. RCAN [13] learns a translation from randomized sim images to a canonical domain, achieving real-world success without real data. GAN-based methods have also shown promising results in translating observations across domains. Sim adaptation techniques aim to align simulation dynamics with reality. AdaptSim [14] meta-learns task-specific simulator parameter adjustments based on real-world performance. Similarly, Scalable Real2Sim [15] uses robotic interaction to infer object parameters and auto-generate simulator assets that can then be used for policy learning. More recent frameworks adopt a real-to-sim-to-real strategy. Torne et al. [16] create real-world digital twins, use inverse policy distillation to initialize sim policies, then fine-tune in simulation for robustness. TRANSIC [17] introduces human-in-the-loop correction to adapt sim policies post-deployment, improving failure recovery. DeepMind’s robot soccer framework [18] combines high-fidelity neural rendering with domain randomization for successful sim-to-real transfer using egocentric vision.
Despite significant progress, most existing methods either aim to learn policies that are robust to sim-to-real discrepancies or rely on fine-tuning in the real world after training in simulation. However, these approaches still suffer from domain shift, misaligned observations, and the need for complex, task-specific simulation modeling. In contrast, our approach seeks to eliminate the sim-to-real gap at its source by keeping the simulator continuously aligned with the real world and treating it as the execution environment throughout data collection, training and deployment. This design decouples the policy from sim-real mismatches and shifts the challenge of domain alignment to the simulator itself—which we update in real time using sensor feedback.
II-C World Models for Robotics
Learned world models allow robots to simulate future outcomes for planning and control, and can act as a natural simulation environment for robotics if trained on specific real-world environments. Traditional models range from 2D black-box predictors to structured simulators with physical priors. UniSim [19] trains a general-purpose, action-conditioned video generator that learns from diverse real-world data, simulating rich observations for both high- and low-level tasks. This approach, however is not well grounded in the 3D physical world and tends to hallucinate beyond a few predicted frames. More recently, 3D physically grounded world models combine particle-based simulation with visual rendering via Gaussian Splatting, enabling real-time tracking and simulated interaction of physical scenes [20, 6] without violating physical and structural constraints. In this work, we explore how they can be utilized within the behavior cloning framework to bridge the advantages of simulation to the real world.
III PRELIMINARIES
III-A Behaviour Cloning
In behavior cloning (BC), the goal is to learn a policy $\pi_{\theta}$ parameterized by $\theta$ that maps an observation $o_{t}$ at time step $t$ to a sequence of future actions $A_{t}=[a_{t},a_{t+1},\ldots,a_{t+H-1}]$ over a horizon $H$. Given a dataset of $N$ expert demonstrations $\mathcal{D}=\{(o^{i},A^{i})\}_{i=1}^{N}$, where $o^{i}=\{o_{1}^{i},\ldots,o_{T}^{i}\}$ is a sequence of observations and $A^{i}=\{a_{1}^{i},\ldots,a_{T}^{i}\}$ is the corresponding sequence of actions, the policy is trained to model the conditional distribution $p(A_{t}\mid o_{t})$. This is typically done by minimizing the behavior cloning loss:
$\mathcal{L}_{\text{BC}}(\pi_{\theta})=\mathbb{E}_{(o,A)\sim\mathcal{D}}\left[\left\|\pi_{\theta}(o)-A\right\|^{2}\right],$ (1)
where the policy’s predicted action sequence is penalized based on its deviation from the expert-provided action sequence.
Modeling temporally extended action sequences rather than single-step actions enables smoother and more consistent control [21, 22]. Our approach adopts Conditional Flow Matching (CFM) [23], a generative modeling method that learns a continuous trajectory distribution over actions. CFM models a velocity field $v^{\theta}(A_{t}^{\tau}\mid o_{t})$ conditioned on $o_{t}$, predicting intermediate velocities along a path $A_{t}^{\tau}$ between a zero trajectory and the ground-truth $A_{t}$. The training objective minimizes the conditional flow-matching loss:
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{\tau,\,p(A_{t}\mid o_{t}),\,q(A_{t}^{\tau}\mid A_{t})}\left\|u(A_{t}^{\tau}\mid A_{t})-v^{\theta}(A_{t}^{\tau}\mid o_{t})\right\|^{2},$ (2)
where $\tau\sim\mathcal{U}(0,1)$ is a uniformly sampled interpolation timestep, $q(A_{t}^{\tau}\mid A_{t})$ defines a reference trajectory (e.g., a Gaussian perturbation around a linear interpolation), and $u(\cdot)$ denotes the corresponding ground-truth velocity.
III-B Problem Formulation
A well-known challenge in behavior cloning is that low training loss $\mathcal{L}_{\text{BC}}$ does not necessarily correlate with high task performance (success rate, reward etc.) during deployment [1, 2, 3]. This disconnect makes regular policy evaluation a crucial component of the training process, as it helps identify optimal checkpoints and avoid both underfitting and overfitting. In simulation environments, such evaluations are feasible due to the ability to parallelize rollouts and perform controlled resets, allowing for frequent and scalable assessment of policy performance. Conversely, real-world evaluations are constrained by practical limitations:
  • Manual Resets: Each evaluation requires manual resetting of the environment, which is labor-intensive and prone to inconsistencies.
  • Sequential Execution: Evaluations must be conducted sequentially, as parallel execution is often impractical, leading to prolonged training cycles.
  • Resource Intensiveness: Continuous human supervision is required to ensure safety and correctness during each trial, further escalating the time and effort involved.
These challenges underscore the need for methodologies that can bridge the gap between the efficiency of simulation-based evaluations and the fidelity of real-world performance, ensuring that policies trained in simulation translate effectively to real-world applications.
Figure 4: Comparison of different representations extracted from the Real-to-Sim framework, contrasted with a real image baseline. (a) A purely state-based representation, utilizing the object’s position and orientation (in quaternion form). (b) A virtual moving camera positioned on the robot’s end effector. (c) A virtual camera placed at the same location as the original camera that observed the scene. (d) One of the real camera images. All representations are augmented with the robot’s state. In the Real-to-Sim framework, the joint states correspond to those of the simulated robot, while in the real image representation, the joint states reflect those of the real robot.
IV METHOD
Our method, real-is-sim, enables seamless integration of parallelisable simulation-based evaluation for real-world robot policies by maintaining a continuously corrected digital twin of the environment. Figure 2(a) summarizes our framework. The policy operates exclusively in the simulator, while the simulator is updated in real time using observations from the real world. This section outlines the full pipeline of real-is-sim, including: (1) our real-time correctable physics simulator built on Embodied Gaussians, (2) the procedure for scene setup and demonstration collection, (3) training policies using Conditional Flow Matching (CFM), and (4) evaluating policies in both offline and online modes.
IV-A Real-Time Correctable Physics Simulator
The cornerstone of real-is-sim is its real-time correctable physics simulator, which maintains synchronization with the real world. Unlike traditional sim-to-real approaches (Figure 2(c)), where the simulator is used only for training and discarded during deployment, our framework keeps the simulator active at all times—during demonstration collection, policy execution, and offline rollouts.
Our implementation leverages the Embodied Gaussian simulator. Embodied Gaussians [6] is a real-time simulator that dynamically aligns with real-world observations through continuous RGB-based correction. It represents the world using a dual-representation which is composed of (1) a set of oriented particles parameterized by their positions, rotations, masses, and radii and (2) a set of 3D Gaussians parameterized by their positions, rotations, scales, color and opacity. The particles are acted upon by a particle-based simulator [24] and represent the physical aspect of the world. In contrast, Gaussians represent the visual aspect of the world because they can be rendered using a differentiable rendering procedure known as Gaussian splatting [25]. The Gaussians are connected to the particles with rigid links. One particle can have multiple Gaussians attached to it. As a particle is moved by the physics system, it simultaneously moves its connected Gaussians as well. The main innovation in Embodied Gaussians is the use of fictitious virtual forces derived from the photometric loss between images received from posed cameras and renders of the scene from the same viewpoints. The displacement of the Gaussians required to minimize the photometric loss is interpreted as a scaled fictitious force that acts on the connected particles. The virtual visual forces are, thus, corrective agents that constantly act on the simulated bodies to align them with what is being observed. The correction loop occurs at 30Hz and allows a physics simulator to constantly synchronize with the real world while always being in a physically feasible state.
Our adaptation (built using [26]) differs from the original framework in two key aspects. First, we forgo shape-matching constraints for deformable objects, instead modeling all entities as rigid bodies composed of spherical collision primitives. Thus, our simulator is a rigid body simulator rather than a particle-based simulator. This simplification improves performance from 30Hz to 60Hz in online mode. Second, we treat the robot as a dynamic entity, simulating mass and inertia for its links and joints. This contrasts with the original kinematic formulation, where robot motion was strictly kinematic. This makes the simulation more stable and it allows the simulated robot to react to the environment.
IV-B Scene Setup
We follow the same procedure outlined in Embodied Gaussians to create the geometries of the objects being manipulated and learn their visual appearance. The result is that each body is represented by a pose $q_{b}\in(\mathbb{R}^{3}\times\mathbb{S}^{3})$ implemented as a position vector and a quaternion, as well as a set of spherical shapes with a fixed radius (empirically tuned) that form the collision geometry of the object. Lastly, the Gaussians forming the visual aspect of the object are learnt with the Gaussian splatting optimization procedure. With the environment constructed and visual appearance learned, we proceed to collect demonstrations using this simulation.
IV-C Demonstration Collection
Demonstrations are collected with Embodied Gaussians in the loop. As the user is collecting demonstrations, the simulator is constantly correcting itself from the observations it gets from the cameras. For every demonstration of a task, we record the full history of simulation states that occurred. Our physics loop runs at 60Hz and we record 60 states every second. The underlying physical model is rich, but the only properties that change over time are the poses of the bodies and the joint configuration of the robot $q_{r}\in\mathbb{R}^{7}$ (for our 7 degree of freedom robot). Since we choose not to allow the visual states to change after they are initialized (for simplicity and to ease the data collection footprint), the state of the Gaussians are only stored at the initial time step. All other physical parameters pertaining to the simulator are also stored once at the initial timestep. A full list of these are documented in [26]. The recorded synchronized simulation trajectories form the dataset used for policy learning.
Figure 5: The success rate of an imitation learning policy as a function of training steps. The plot compares policy evaluation in the offline mode of real-is-sim with the online mode of real-is-sim. The high correlation between the two curves demonstrates that the (significantly faster) offline evaluation mode is a reliable proxy for real-world testing, offering a practical alternative for policy assessment. Error bars are 95% confidence intervals.
IV-D Policy Training
Our policy outputs two components: $q_{d}\in\mathbb{R}^{7}$, the desired joint configuration of the robot, and a scalar progress value $p\in[0,1]$, which indicates the relative completion (progress) of the task. Progress is a heuristic with a ground-truth label that is calculated as the relative time elapsed in the demonstration. It assumes that every demonstration ends at the target position. This was shown to work well as a heuristic in [27]. We use it in our work to evaluate the trained policy without needing to engineer a reward. Our policy is conditioned on the current joint configuration of the simulated robot. In the experiment section, state-based policies are also conditioned on an object’s pose encoded as a position and quaternion (a quaternion is used for simplicity because the pose is stored natively in that form and because it’s double cover property did not prove to be a problem for our task). For image-based policies we use a ResNet to encode an image into a 64 dimensional vector which we condition on. For image-based policies, we found that random crops were crucial for successful learning. Embodied Gaussians allows us to render images from any viewpoint and create virtual cameras in the setup. The representations used to condition the policy—including object pose and rendered images—are explored further in Section V-D. Our policies output 32 actions which are meant to be executed at 30Hz. During evaluation, we execute all 32 actions before resampling.
IV-E Offline Policy Evaluation
During offline policy evaluation, we create 20 parallel environments as shown in Figure 3(b) and run the policies for a predetermined time (90 seconds). We regard any policy that has not reached a progress of above 0.9 as a failure. For state-based policies with 20 environments, we can run the simulation with policy evaluation at 8x realtime on an NVIDIA 6000Ada. With one virtual camera rendering the scene in every environment, the speed is 5x realtime.
V EXPERIMENTS
Figure 6: Success rates of different demonstration collection strategies. A dataset consisting of 30 real demonstrations achieves a success rate of approximately 58%. Since Real-to-Sim can also operate offline, the demonstrations can be augmented with additional simulated data. A combined dataset of 60 demonstrations (30 real and 30 simulated) achieves an 80% success rate. In comparison, a dataset of 60 real demonstrations alone achieves a slightly lower success rate of 72%. Error bars are 95% confidence intervals.
Figure 7: Success rates of policies using different representations to condition the flow matching policy. The baseline image-based policy (not using real-is-sim) uses a static camera and has a success rate of 63%. The state-based policy, which uses the simulator’s object pose achieves a higher success rate of 76%. Lastly, policies trained on the rendered virtual cameras (one statically mounted and another mounted on the gripper) achieve 73% and 82% respectively. This shows that a gripper-mounted camera is likely a better representation for the PushT task. Error bars are 95% confidence intervals.
V-A Experimental Setup
We evaluate our framework on the PushT task: a long horizon task where a robot uses a pusher to position a T-shaped object into a predefined central location. The T-Block’s geometry (represented as a set of spherical shapes attached to a single body frame) is automatically generated using the method from Embodied Gaussians [6]. We use three RealSense D455 cameras running at 90 fps with a resolution of $420\times 270$. The robot mesh is known, and the ground plane is estimated from the camera-generated pointclouds. Gaussians attached to the robot, table, and T-Block are learned once using the Gaussian Splatting procedure outline in [6] and reused across experiments.
For state-based policies, we use the pose of the object encoded as a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\in\mathbb{SO}(3)$, where $\mathbf{t}=(x,y,z)$ and $\mathbf{q}=(q_{x},q_{y},q_{z},q_{w})$. For image-based policies, we use a 64 dimensional feature vector output by a Resnet [22]. The images input to the Resnet are rendered from the simulation using Gaussian Splatting from virtual cameras. These cameras can be placed anywhere. In our experiments, we use a static camera placed in the same location as the real camera and a virtual gripper camera which we mount to the simulated robot’s end-effector (these are shown in Figure 7).
To measure policy success on the PushT task, we define 20 in-distribution starting poses on the table that were not part of the demonstration set. In real-world testing, success is determined by whether the policy drives the T-Block to the target location from these poses. Each pose is tested three times, resulting in 60 measurements per success rate evaluation. Error bars in figures represent 95% confidence intervals.
V-B Offline Evaluation
In this experiment, we demonstrate how real-is-sim’s offline mode facilitates policy training and evaluation. A key challenge in imitation learning lies in selecting optimal checkpoints for deployment. This difficulty arises because the loss used in behavior cloning does not reliably indicate policy performance, and real-world robot evaluations require substantial time and resources. Consequently, researchers often default to using the final training checkpoints despite it being suboptimal.
We demonstrate that real-is-sim’s offline mode enables efficient checkpoint selection through rapid evaluation. For a state-based policy trained on 30 demonstrations, we implement a flow-matching policy trained for 50,000 steps (Figure 5). We then compare two evaluation methodologies: (i) offline rollouts using real-is-sim and (ii) real-world online evaluations. Our offline approach initializes environments with identical test cases to online evaluations. The offline evaluation calculates success rates through the policy-reported progress. While we use this metric to maintain generality, practitioners could implement more sophisticated reward functions for their specific tasks.
We selected four representative checkpoints (10k, 15k, 25k, 50k) for online evaluation (each requires around 1 hour to perform) while evaluating more comprehensively through offline mode. Figure 5 reveals strong correlation between evaluation curves, demonstrating that faster offline evaluations (minutes vs. hours) effectively approximate real-world results. Notably, later checkpoints do not consistently outperform earlier ones, suggesting practitioners could use offline evaluation for either early stopping or optimal checkpoint selection. While this experiment uses identical initial states for both evaluation modes to establish correlation, real applications could benefit from randomized initial states during training-phase evaluations.
V-C Offline Data Collection and Augmentation
In this experiment, we demonstrate how real-is-sim’s offline mode can be used to to collect additional demonstrations to increase task performance (Figure 6). We train a state-based policy using 30 real-world demonstrations collected in online mode - which achieves a success rate of around 57%. We also augment the real-world demonstrations using the simulator only (offline mode). To collect the additional demonstrations, we deploy this policy in offline mode across multiple parallel environments and identify failure states where the policy falls into a local minima. These failure states are then presented to the user to provide an additional 30 demonstrations specifically solving the task under these challenging configurations. By combining the original 30 real-world demonstrations with the 30 new simulated demonstrations, we retrain the policy and observe a significant improvement to 80% success rate.
Lastly as a baseline, we also collect 30 demonstrations in simulation only. The resulting sim-only policy attains a similar success rate during deployment. The success rate of the offline only policy is comparable to the policy trained on real demonstrations. This is largely owing to the kinematic nature of the PushT task which means that successfully policy rollouts are mostly dependent on poses and not dynamic effects. This is precisely what Embodied Gaussians can easily correct upon deployment.
This experiment highlights a straightforward yet powerful application of real-is-sim’s offline mode. The cost of augmenting the data used for policy training using real-is-sim’s simulator is much lower than doing so in the real world. As a comparison, we also train a policy using 60 real-world demonstrations collected without augmentation. Interestingly, this approach yields a similar performance improvement as the augmented dataset, underscoring the effectiveness of real-is-sim’s offline mode in efficiently enhancing policy training.
V-D Representation Flexibility
Another major benefit of real-is-sim’s offline mode is its flexibility in extracting diverse representations from the Embodied Gaussians simulation. For instance, it enables access to privileged information—such as precise object poses—that would be difficult to obtain in the real world. Additionally, virtual cameras can be freely positioned within the system, facilitating visuomotor policy learning from multiple perspectives. In our experiments, we demonstrate that policies leveraging different representations—including state-based object poses, a gripper-mounted virtual camera, and a static virtual camera—all successfully solve the task (Figure 7, 4). Notably, these approaches achieve performance comparable to, and in some cases surpassing, the baseline policy that relies solely on real-world images and robot actions.
Interestingly, the policy using the gripper-mounted virtual camera achieves the highest performance (82%) among all representations. This policy also exhibits behaviors such as as actively searching for the T-block when it leaves the camera’s field of view (see videos on our website). These findings align with prior work highlighting the advantages of wrist-mounted cameras over static ones in certain scenarios [28]. Within the real-to-sim framework, we can systematically evaluate different virtual camera configurations [29] and their impact on policy learning—all using the same set of demonstrations. This flexibility allows us to identify the optimal representation for a given manipulation task in offline mode and seamlessly deploy it in online execution.
VI Limitations
While our framework demonstrates promising results, its applicability is inherently constrained by the capabilities of the underlying physical simulator. Scenarios that deviate significantly from the simulator’s modeling assumptions pose challenges to both model accuracy and corrective mechanisms. To address these constraints, future improvements could focus on enhancing simulator fidelity through adaptive parameter estimation (e.g., real-time optimization of physical properties like friction or mass, as explored in works such as [20] and  [30]). While this work focused on the PushT task to validate core contributions, the methodology is extensible to broader manipulation domains. Enabling applications such as multi-object rearrangement, cloth manipulation, or non-prehensile pushing will require tailoring the simulator’s dynamics model and corrective policies to handle domain-specific interactions.
VII Future Work
While the always-in-the-loop simulator already corrects itself using observations, its accuracy could be further enhanced by backpropagating into physics parameters or learning a residual physics network, reducing reliance on explicit corrections by making the simulator’s predictions more physically faithful. A more accurate simulator would also enable reinforcement-learning-based fine-tuning of imitation-learned policies. Additionally, this work focused on offline policy evaluation, but future extensions could incorporate online planning during deployment for real-time adaptability. Finally, Embodied Gaussians could be extended to encode higher-level features such as DINO embeddings [31] or other semantic descriptors. This would enrich policy training and further make use of the simulator’s role as a mediator between the real world and the policy.
VIII CONCLUSIONS
We introduced real-is-sim, a unified framework for policy training, evaluation, and deployment that maintains a continuously corrected simulation throughout the entire pipeline. By treating simulation as a persistent, real-time interface, enabled by Embodied Gaussians, we eliminate the disconnect between training and deployment domains. This reframes policy evaluation as a simulation-native process, even in real-world settings. Our results on the PushT task show strong alignment between offline and real-world performance, demonstrating that real-is-sim enables scalable, efficient evaluation without sacrificing fidelity. We hope this framework serves as a foundation for future real-world learning algorithms that leverage this tight coupling between simulation and reality.

### v2

@@
-Real-is-Sim: Bridging the Sim-to-Real Gap with a Dynamic Digital Twin for Real-World Robot Policy Evaluation
+Real-is-Sim: Bridging the Sim-to-Real Gap with a Dynamic Digital Twin
 Abstract
-Recent advancements in behavior cloning have enabled robots to perform complex manipulation tasks.
-However, accurately assessing training performance remains challenging, particularly for real-world applications, as behavior cloning losses often correlate poorly with actual task success.
-Consequently, researchers resort to success rate metrics derived from costly and time-consuming real-world evaluations, making the identification of optimal policies and detection of overfitting or underfitting impractical.
-To address these issues, we propose real-is-sim, a novel behavior cloning framework that incorporates a dynamic digital twin (based on Embodied Gaussians) throughout the entire policy development pipeline: data collection, training, and deployment.
-By continuously aligning the simulated world with the physical world, demonstrations can be collected in the real world with states extracted from the simulator.
-The simulator enables flexible state representations by rendering image inputs from any viewpoint or extracting low-level state information from objects embodied within the scene.
-During training, policies can be directly evaluated within the simulator in an offline and highly parallelizable manner.
-Finally, during deployment, policies are run within the simulator where the real robot directly tracks the simulated robot’s joints, effectively decoupling policy execution from real hardware and mitigating traditional domain-transfer challenges.
-We validate real-is-sim on the PushT manipulation task, demonstrating strong correlation between success rates obtained in the simulator and real-world evaluations.
-Videos of our system can be found at https://realissim.rai-inst.com.
-I INTRODUCTION
+We introduce real-is-sim, a new approach to integrating simulation into behavior cloning pipelines.
+In contrast to real-only methods, which lack the ability to safely test policies before deployment, and sim-to-real methods, which require complex adaptation to cross the sim-to-real gap, our framework allows policies to seamlessly switch between running on real hardware and running in parallelized virtual environments.
+At the center of real-is-sim is a dynamic digital twin, powered by the Embodied Gaussian simulator, that synchronizes with the real world at 60Hz.
+This twin acts as a mediator between the behavior cloning policy and the real robot.
+Policies are trained using representations derived from simulator states and always act on the simulated robot, never the real one.
+During deployment, the real robot simply follows the simulated robot’s joint states, and the simulation is continuously corrected with real world measurements.
+This setup, where the simulator drives all policy execution and maintains real-time synchronization with the physical world, shifts the responsibility of crossing the sim-to-real gap to the digital twin’s synchronization mechanisms, instead of the policy itself.
+We demonstrate real-is-sim on a long-horizon manipulation task (PushT), showing that virtual evaluations are consistent with real-world results.
+We further show how real-world data can be augmented with virtual rollouts and compare to policies trained on different representations derived from the simulator state including object poses and rendered images from both static and robot-mounted cameras.
+Our results highlight the flexibility of the real-is-sim framework across training, evaluation, and deployment stages.
+Videos available at https://real-is-sim.github.io.
+1 INTRODUCTION
+Simulation environments are a powerful tool for developing robotic manipulation policies.
+They offer key advantages such as continuous performance monitoring, large-scale parallel rollouts, full access to state information, and precise control over resets and environmental variation.
+These features make policy learning in simulation significantly more practical to do efficiently.
+However, translating these advantages to the real world is challenging.
+The main difficulty lies in ensuring that a policy behaves the same way in both simulation and reality.
+This requires two things: aligning the input distributions the policy sees in both domains, and ensuring that both the simulated and real robots respond similarly to the same policy outputs.
+Achieving both reliably is difficult.
 Figure 1: The real-is-sim framework, illustrating the information flow between its components.
-A policy trained in a correctable physics simulator (leveraging Embodied Gaussians) controls a simulated robot.
+A policy trained in a physics simulator that can be synchronized with the real world [1] controls a simulated robot.
 Real-world observations continuously update the simulator, maintaining its state close to ground truth.
 The real robot then mirrors the simulated robot’s joint positions.
-This approach shifts the sim-to-real gap challenge from the policy to the adaptable physics simulator.
-Simulation environments offer a highly favorable setting for developing and evaluating robotic manipulation policies.
-They support continuous monitoring of task performance during training, enable large-scale parallel rollouts, and allow for precise control over resets and environment variation.
-Crucially, simulation also affords flexible and often privileged access to state information, such as exact object poses or custom visual observations from arbitrary viewpoints, that can simplify policy learning.
-These advantages make behavior cloning particularly effective in simulation.
-However, transferring these benefits to the real world remains challenging.
-Real-world policy evaluation is slow, labor-intensive, and limited in scale, often relying on training loss as a proxy for success despite its poor correlation with actual task performance [1, 2, 3].
-This gap between simulation and real-world applicability motivates the core question driving our work: How can we retain the strengths of simulation for behavior cloning while operating in the real world?
-Despite the clear advantages of simulation, integrating these benefits into real-world policy development remains difficult.
-Broadly, two paradigms exist.
-The first, real-only training, relies exclusively on physical hardware [4].
-While this offers immediate applicability, since policies are trained and deployed in the same environment, it comes with significant limitations: (i) it prevents offline evaluation, which is essential for diagnosing failure modes and tracking performance, and (ii) it restricts large-scale parallelization, which is useful for exploration, hyperparameter tuning, and proactive planning.
-The second paradigm, sim-to-real, aims to overcome these constraints by training in simulation and transferring policies to the real world.
-This allows for extensive offline evaluation and experimentation at scale [5].
-However, it introduces the well-known sim-to-real gap: discrepancies in observations and dynamics between simulation and reality that can degrade policy performance.
-Bridging this gap often requires extensive, task-specific tuning of either the simulator or the policy, efforts that are difficult to generalize or scale.
-Current approaches treat simulation and reality as separate domains, creating a mismatch between the observational inputs a learned manipulation policy receives during training and deployment.
-This discrepancy also necessitates ensuring that the simulated robot behaves identically to its real-world counterpart, making policy transfer from simulation to reality challenging.
-We propose real-is-sim, a paradigm shift where a correctable simulator remains always-in-the-loop, even during real-world deployment.
-A correctable simulator synchronizes with the physical world using sensor data (in our case, RGB images), creating a live, dynamic digital twin of the robot’s environment.
-The first simulator to achieve this generally and in real-time is Embodied Gaussians [6].
-Embodied Gaussians combines the use of Gaussian splatting and a particle-based physics system to represent both the visual and physical aspects of the world.
-It allows, through its differentiable rendering process, the correction of the simulator state from observations received from cameras.
-In the real-is-sim framework, the always-in-the-loop simulator acts as a mediator between the physical world and the policy.
-Policies are trained exclusively on representations derived from the simulator’s state, effectively transforming real-world interactions into simulated states.
-This approach collapses the traditional divide between training and deployment domains.
-Real-is-sim collects synchronized simulation states during real-world interactions, which serve as demonstration trajectories for policy training.
-Policies are trained on arbitrary representations extracted from these simulated states (which may include object poses, robot configuration, and/or rendered images from fixed or robot-mounted viewpoints).
-The policies output desired joint-space actions that control the simulated robot.
-Critically, this process is identical in both online and offline operational modes.
-In offline mode, the simulator operates without real-world coupling, executing the policy in a purely virtual environment.
-In online mode, the real robot acts as a follower and tracks the simulated robot’s joint positions, while real-world sensor data (RGB cameras) continuously correct the simulator’s state.
-This design (illustrated in Figure 1) ensures the policy remains fully decoupled from the physical hardware — it interacts solely with the simulated robot and environment in all modes.
-As such it allows the policies to always operate on in-distribution simulator states, even when deployed in the real world.
-By maintaining a persistent, real-time digital twin of the physical environment, real-is-sim unifies simulation and reality into a single, coherent interface for policy development.
-This operational symmetry across both offline and online modes unlocks several key benefits: consistent policy behavior between simulation and deployment, scalable and safe offline evaluation, and flexible, perception-agnostic representations for policy learning.
-In this work, we show how real-is-sim fundamentally rethinks the role of simulation, not just as a training tool, but as a medium for end-to-end policy evaluation and deployment.
-We validate our framework on the PushT manipulation task, demonstrating that offline simulation evaluations closely predict real-world performance.
-Real-is-sim offers a scalable path forward for efficient training, reliable assessment, and robust deployment of robotic policies in the real world.
-Figure 2: Comparison of three paradigms: Real-is-Sim, Real-Only, and Sim-to-Real, highlighting their online (real-world interaction) and offline (simulation-based) capabilities.
-Real-is-Sim offers a unified framework where the offline mode is a simplified version of the online, lacking only real-time correction.
-This ensures seamless transferability.
-Real-Only is confined to real-world execution.
-Sim-to-Real struggles with distribution mismatch and dynamics discrepancies, making successful transfer uncertain.
-In the Real-is-Sim paradigm, successful transfer between online and offline modes is mostly dependent on whether environment synchronization was successful during the demo collection phase.
-II RELATED WORK
-II-A Real World Policy Evaluation
-Evaluating robotic policies directly in the real world remains the most reliable method for assessing performance in practical tasks [4]; however, it is often labor-intensive, time-consuming, and requires substantial human oversight [1].
-In many experimental pipelines, researchers must manually reset the environment, initiate policy execution, and supervise task completion.
-For example, in the benchmark introduced by Cheng et al.
-[7], the authors report manually aligning a T-shaped object to match one of 20 predefined start poses for each policy rollout.
-This reset process involved visual overlays to ensure pose accuracy and was repeated across all baselines.
-This process is not parallelizable and poses a scalability bottleneck when evaluating across multiple tasks, policies, and environment variations.
-To mitigate these challenges, a growing body of work has turned to evaluating real-world policies in simulation environments, even when those policies are trained on real data [8, 9, 10, 11].
-This strategy offers safety, repeatability, and especially scalability, enabling parallelized policy assessment across diverse conditions.
-Prior works have spent large amounts of effort trying to improve the correlation between simulated and real world performance by improving visual appearance [8, 11], refining object models [8, 10], modeling control latency [9], and tracking cross-corelation between object classes [11].
-In contrast to these evaluation-centric approaches, our framework incorporates the simulator as an integral component throughout the entire policy development pipeline: training, evaluation, and deployment.
-Rather than constructing simulation for post hoc policy assessment, we instead continuously synchronize the simulation with real-world sensor data during execution.
-This enables policies trained entirely in simulation to be deployed directly on real robots without additional fine-tuning or adaptation.
-By maintaining real-time alignment between the simulated and physical environments, our method ensures that the policy operates within a unified and consistent observation and control space.
-This paradigm shift facilitates seamless transfer between simulation and reality, enabling scalable evaluation and robust deployment under a single, coherent framework.
-Figure 3: The two modes of the Real-is-Sim framework: online and offline.
-In online mode, the real robot operates alongside the correctable simulator (Embodied Gaussians).
-Synchronization is maintained through fictitious visual forces derived from the rendering loss between observed and real images, as well as by the real robot mimicking the simulated robot’s state.
-In offline mode, parallel environments execute the learned policy and evaluate its success rate.
-The simulator can run up to 7x faster than real time with policy sampling in the loop.
-II-B Bridging the Sim-to-Real Gap for Robotic Manipulation
-As discussed above, apart from direct real world evaluation [4], simulation of the corresponding real world environment offers a promising alternative.
-A major challenge lies in the discrepancy between the simulated states and reality particularly.
-To bridge this gap, numerous strategies have been proposed.
-Domain randomization injects variability into the simulation (e.g., textures, lighting, physics) to force robustness.
-Early work by Tobin et al.
-[12] demonstrated zero-shot sim-to-real transfer using randomized visuals.
-While powerful, this approach often requires excessive variability and long training times.
-Domain adaptation methods attempt to align visual or latent representations between sim and real.
-RCAN [13] learns a translation from randomized sim images to a canonical domain, achieving real-world success without real data.
-GAN-based methods have also shown promising results in translating observations across domains.
-Sim adaptation techniques aim to align simulation dynamics with reality.
-AdaptSim [14] meta-learns task-specific simulator parameter adjustments based on real-world performance.
-Similarly, Scalable Real2Sim [15] uses robotic interaction to infer object parameters and auto-generate simulator assets that can then be used for policy learning.
-More recent frameworks adopt a real-to-sim-to-real strategy.
-Torne et al.
-[16] create real-world digital twins, use inverse policy distillation to initialize sim policies, then fine-tune in simulation for robustness.
-TRANSIC [17] introduces human-in-the-loop correction to adapt sim policies post-deployment, improving failure recovery.
-DeepMind’s robot soccer framework [18] combines high-fidelity neural rendering with domain randomization for successful sim-to-real transfer using egocentric vision.
-Despite significant progress, most existing methods either aim to learn policies that are robust to sim-to-real discrepancies or rely on fine-tuning in the real world after training in simulation.
-However, these approaches still suffer from domain shift, misaligned observations, and the need for complex, task-specific simulation modeling.
-In contrast, our approach seeks to eliminate the sim-to-real gap at its source by keeping the simulator continuously aligned with the real world and treating it as the execution environment throughout data collection, training and deployment.
-This design decouples the policy from sim-real mismatches and shifts the challenge of domain alignment to the simulator itself—which we update in real time using sensor feedback.
-II-C World Models for Robotics
-Learned world models allow robots to simulate future outcomes for planning and control, and can act as a natural simulation environment for robotics if trained on specific real-world environments.
-Traditional models range from 2D black-box predictors to structured simulators with physical priors.
-UniSim [19] trains a general-purpose, action-conditioned video generator that learns from diverse real-world data, simulating rich observations for both high- and low-level tasks.
-This approach, however is not well grounded in the 3D physical world and tends to hallucinate beyond a few predicted frames.
-More recently, 3D physically grounded world models combine particle-based simulation with visual rendering via Gaussian Splatting, enabling real-time tracking and simulated interaction of physical scenes [20, 6] without violating physical and structural constraints.
-In this work, we explore how they can be utilized within the behavior cloning framework to bridge the advantages of simulation to the real world.
-III PRELIMINARIES
-III-A Behaviour Cloning
+This approach shifts the sim-to-real gap challenge from the policy to the physics simulator.
+To overcome this limitation, we introduce real-is-sim, a new paradigm in which a simulator that continuously synchronizes with the real world is the sole interface to a behavior cloning policy and is ever-present; even during real-world execution.
+This paradigm is powered by Embodied Gaussians [1], the first simulator to achieve real-time synchronization with physical environments in a generalizable way, through the integration of Gaussian splatting [2] and particle-based physics [3].
+This simulator enables the creation of a dynamic digital twin of the physical system that can be updated in real time via sensor feedback.
+Policies are trained on representations derived from the simulator’s state, effectively reframing real-world interactions as simulated experiences and unifying training and deployment.
+Real-is-sim functions by collecting synchronized simulation states during real-world interactions, which serve as demonstration trajectories for policy training.
+We avoid sim-only demonstrations to ensure all trajectories reflect motions the real robot can reliably execute whilst keeping the simulation synchronized with the real world.
+These policies operate on simulator-derived representations and produce joint-space actions for the simulated robot.
+This process is consistent in the two main operating modes: (i) when deployed on a real robot, and (ii) when evaluating in virtual environments.
+When connected to a real robot, the physical robot mirrors the simulated robot’s configuration, while sensor data continuously updates the state of the objects in the simulator.
+This architecture decouples the policy from the physical hardware, ensuring the policy always operates on simulator states drawn from a similar distribution as the training data.
+The real-is-sim paradigm enables seamless switching between real-world deployment and virtual evaluation, as the only distinction between the two modes is the activation or deactivation of the real robot’s follower behavior.
+By maintaining a real-time digital twin, the real-is-sim framework unifies simulation and reality, enabling consistent policy execution, scalable offline evaluation, and flexible representation learning.
+Our experiments on the PushT task demonstrate three key applications of real-is-sim: (i) enabling rapid policy evaluation in a virtual environment that correlates strongly with real-world performance, (ii) supporting data augmentation by intermixing real and simulated data to improve performance, and (iii) identifying optimal simulation-derived representations to maximize task effectiveness.
+2 PRELIMINARIES
+2.1 Digital Twin
+The real-is-sim framework is built around a simulator that stays synchronized with the real world.
+This is achieved using Embodied Gaussians [1], a simulator that continuously corrects itself using RGB-based feedback from real-world observations.
+The system represents the environment using two tightly coupled components:
+  1. 1.
+    Oriented particles, which capture physical properties such as position, rotation, mass, and radius. These are simulated using a particle-based physics system [3].
+  2. 2.
+    3D Gaussians, which model visual appearance through parameters such as position, rotation, scale, color, and opacity. These are rendered using Gaussian splatting [2].
+Figure 2: Figure adapted from [1].
+It illustrates how Embodied Gaussians represents the world using physical particles and visual Gaussians.
+The Gaussians are corrected using real RGB images, and in turn exert fictitious visual forces on the particles to align the simulation with the real world.
+The adaptation highlights how this mechanism integrates into the real-is-sim framework.
+Each particle can be rigidly linked to multiple Gaussians, allowing the physical and visual representations to move together.
+As the physics system updates the particles, the attached Gaussians follow accordingly.
+The key innovation in Embodied Gaussians is its correction mechanism based on photometric error.
+By comparing real camera images with rendered views, the system computes fictitious visual forces that act on the particles.
+These forces continuously adjust the simulation to stay aligned with reality.
+This correction loop runs at 60Hz, ensuring the simulator remains both physically consistent and visually accurate.
+An illustration of the corrective mechanism is provided in Figure 2.
+Additional details are available in [1].
+2.2 Behaviour Cloning
 In behavior cloning (BC), the goal is to learn a policy $\pi_{\theta}$ parameterized by $\theta$ that maps an observation $o_{t}$ at time step $t$ to a sequence of future actions $A_{t}=[a_{t},a_{t+1},\ldots,a_{t+H-1}]$ over a horizon $H$.
@@
 The training objective minimizes the conditional flow-matching loss:
-$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{\tau,\,p(A_{t}\mid o_{t}),\,q(A_{t}^{\tau}\mid A_{t})}\left\|u(A_{t}^{\tau}\mid A_{t})-v^{\theta}(A_{t}^{\tau}\mid o_{t})\right\|^{2},$ (2)
-where $\tau\sim\mathcal{U}(0,1)$ is a uniformly sampled interpolation timestep, $q(A_{t}^{\tau}\mid A_{t})$ defines a reference trajectory (e.g., a Gaussian perturbation around a linear interpolation), and $u(\cdot)$ denotes the corresponding ground-truth velocity.
-III-B Problem Formulation
-A well-known challenge in behavior cloning is that low training loss $\mathcal{L}_{\text{BC}}$ does not necessarily correlate with high task performance (success rate, reward etc.) during deployment [1, 2, 3].
-This disconnect makes regular policy evaluation a crucial component of the training process, as it helps identify optimal checkpoints and avoid both underfitting and overfitting.
-In simulation environments, such evaluations are feasible due to the ability to parallelize rollouts and perform controlled resets, allowing for frequent and scalable assessment of policy performance.
-Conversely, real-world evaluations are constrained by practical limitations:
-  • Manual Resets: Each evaluation requires manual resetting of the environment, which is labor-intensive and prone to inconsistencies.
-  • Sequential Execution: Evaluations must be conducted sequentially, as parallel execution is often impractical, leading to prolonged training cycles.
-  • Resource Intensiveness: Continuous human supervision is required to ensure safety and correctness during each trial, further escalating the time and effort involved.
-These challenges underscore the need for methodologies that can bridge the gap between the efficiency of simulation-based evaluations and the fidelity of real-world performance, ensuring that policies trained in simulation translate effectively to real-world applications.
-Figure 4: Comparison of different representations extracted from the Real-to-Sim framework, contrasted with a real image baseline.
-(a) A purely state-based representation, utilizing the object’s position and orientation (in quaternion form).
-(b) A virtual moving camera positioned on the robot’s end effector.
-(c) A virtual camera placed at the same location as the original camera that observed the scene.
-(d) One of the real camera images.
-All representations are augmented with the robot’s state.
-In the Real-to-Sim framework, the joint states correspond to those of the simulated robot, while in the real image representation, the joint states reflect those of the real robot.
-IV METHOD
-Our method, real-is-sim, enables seamless integration of parallelisable simulation-based evaluation for real-world robot policies by maintaining a continuously corrected digital twin of the environment.
-Figure 2(a) summarizes our framework.
-The policy operates exclusively in the simulator, while the simulator is updated in real time using observations from the real world.
-This section outlines the full pipeline of real-is-sim, including: (1) our real-time correctable physics simulator built on Embodied Gaussians, (2) the procedure for scene setup and demonstration collection, (3) training policies using Conditional Flow Matching (CFM), and (4) evaluating policies in both offline and online modes.
-IV-A Real-Time Correctable Physics Simulator
-The cornerstone of real-is-sim is its real-time correctable physics simulator, which maintains synchronization with the real world.
-Unlike traditional sim-to-real approaches (Figure 2(c)), where the simulator is used only for training and discarded during deployment, our framework keeps the simulator active at all times—during demonstration collection, policy execution, and offline rollouts.
-Our implementation leverages the Embodied Gaussian simulator.
-Embodied Gaussians [6] is a real-time simulator that dynamically aligns with real-world observations through continuous RGB-based correction.
-It represents the world using a dual-representation which is composed of (1) a set of oriented particles parameterized by their positions, rotations, masses, and radii and (2) a set of 3D Gaussians parameterized by their positions, rotations, scales, color and opacity.
-The particles are acted upon by a particle-based simulator [24] and represent the physical aspect of the world.
-In contrast, Gaussians represent the visual aspect of the world because they can be rendered using a differentiable rendering procedure known as Gaussian splatting [25].
-The Gaussians are connected to the particles with rigid links.
-One particle can have multiple Gaussians attached to it.
-As a particle is moved by the physics system, it simultaneously moves its connected Gaussians as well.
-The main innovation in Embodied Gaussians is the use of fictitious virtual forces derived from the photometric loss between images received from posed cameras and renders of the scene from the same viewpoints.
-The displacement of the Gaussians required to minimize the photometric loss is interpreted as a scaled fictitious force that acts on the connected particles.
-The virtual visual forces are, thus, corrective agents that constantly act on the simulated bodies to align them with what is being observed.
-The correction loop occurs at 30Hz and allows a physics simulator to constantly synchronize with the real world while always being in a physically feasible state.
-Our adaptation (built using [26]) differs from the original framework in two key aspects.
-First, we forgo shape-matching constraints for deformable objects, instead modeling all entities as rigid bodies composed of spherical collision primitives.
-Thus, our simulator is a rigid body simulator rather than a particle-based simulator.
-This simplification improves performance from 30Hz to 60Hz in online mode.
-Second, we treat the robot as a dynamic entity, simulating mass and inertia for its links and joints.
-This contrasts with the original kinematic formulation, where robot motion was strictly kinematic.
-This makes the simulation more stable and it allows the simulated robot to react to the environment.
-IV-B Scene Setup
-We follow the same procedure outlined in Embodied Gaussians to create the geometries of the objects being manipulated and learn their visual appearance.
-The result is that each body is represented by a pose $q_{b}\in(\mathbb{R}^{3}\times\mathbb{S}^{3})$ implemented as a position vector and a quaternion, as well as a set of spherical shapes with a fixed radius (empirically tuned) that form the collision geometry of the object.
-Lastly, the Gaussians forming the visual aspect of the object are learnt with the Gaussian splatting optimization procedure.
-With the environment constructed and visual appearance learned, we proceed to collect demonstrations using this simulation.
-IV-C Demonstration Collection
-Demonstrations are collected with Embodied Gaussians in the loop.
-As the user is collecting demonstrations, the simulator is constantly correcting itself from the observations it gets from the cameras.
-For every demonstration of a task, we record the full history of simulation states that occurred.
-Our physics loop runs at 60Hz and we record 60 states every second.
-The underlying physical model is rich, but the only properties that change over time are the poses of the bodies and the joint configuration of the robot $q_{r}\in\mathbb{R}^{7}$ (for our 7 degree of freedom robot).
-Since we choose not to allow the visual states to change after they are initialized (for simplicity and to ease the data collection footprint), the state of the Gaussians are only stored at the initial time step.
-All other physical parameters pertaining to the simulator are also stored once at the initial timestep.
-A full list of these are documented in [26].
-The recorded synchronized simulation trajectories form the dataset used for policy learning.
-Figure 5: The success rate of an imitation learning policy as a function of training steps.
-The plot compares policy evaluation in the offline mode of real-is-sim with the online mode of real-is-sim.
-The high correlation between the two curves demonstrates that the (significantly faster) offline evaluation mode is a reliable proxy for real-world testing, offering a practical alternative for policy assessment.
-Error bars are 95% confidence intervals.
-IV-D Policy Training
-Our policy outputs two components: $q_{d}\in\mathbb{R}^{7}$, the desired joint configuration of the robot, and a scalar progress value $p\in[0,1]$, which indicates the relative completion (progress) of the task.
-Progress is a heuristic with a ground-truth label that is calculated as the relative time elapsed in the demonstration.
-It assumes that every demonstration ends at the target position.
-This was shown to work well as a heuristic in [27].
-We use it in our work to evaluate the trained policy without needing to engineer a reward.
-Our policy is conditioned on the current joint configuration of the simulated robot.
-In the experiment section, state-based policies are also conditioned on an object’s pose encoded as a position and quaternion (a quaternion is used for simplicity because the pose is stored natively in that form and because it’s double cover property did not prove to be a problem for our task).
-For image-based policies we use a ResNet to encode an image into a 64 dimensional vector which we condition on.
-For image-based policies, we found that random crops were crucial for successful learning.
-Embodied Gaussians allows us to render images from any viewpoint and create virtual cameras in the setup.
-The representations used to condition the policy—including object pose and rendered images—are explored further in Section V-D.
-Our policies output 32 actions which are meant to be executed at 30Hz.
-During evaluation, we execute all 32 actions before resampling.
-IV-E Offline Policy Evaluation
-During offline policy evaluation, we create 20 parallel environments as shown in Figure 3(b) and run the policies for a predetermined time (90 seconds).
-We regard any policy that has not reached a progress of above 0.9 as a failure.
-For state-based policies with 20 environments, we can run the simulation with policy evaluation at 8x realtime on an NVIDIA 6000Ada.
-With one virtual camera rendering the scene in every environment, the speed is 5x realtime.
-V EXPERIMENTS
-Figure 6: Success rates of different demonstration collection strategies.
-A dataset consisting of 30 real demonstrations achieves a success rate of approximately 58%.
-Since Real-to-Sim can also operate offline, the demonstrations can be augmented with additional simulated data.
-A combined dataset of 60 demonstrations (30 real and 30 simulated) achieves an 80% success rate.
-In comparison, a dataset of 60 real demonstrations alone achieves a slightly lower success rate of 72%.
-Error bars are 95% confidence intervals.
-Figure 7: Success rates of policies using different representations to condition the flow matching policy.
-The baseline image-based policy (not using real-is-sim) uses a static camera and has a success rate of 63%.
-The state-based policy, which uses the simulator’s object pose achieves a higher success rate of 76%.
-Lastly, policies trained on the rendered virtual cameras (one statically mounted and another mounted on the gripper) achieve 73% and 82% respectively.
-This shows that a gripper-mounted camera is likely a better representation for the PushT task.
-Error bars are 95% confidence intervals.
-V-A Experimental Setup
-We evaluate our framework on the PushT task: a long horizon task where a robot uses a pusher to position a T-shaped object into a predefined central location.
+$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{\tau,\,p(A_{t}\mid o_{t}),\,q_{r}(A_{t}^{\tau}\mid A_{t})}\left\|u(A_{t}^{\tau}\mid A_{t})-v^{\theta}(A_{t}^{\tau}\mid o_{t})\right\|^{2},$ (2)
+where $\tau\sim\mathcal{U}(0,1)$ is a uniformly sampled interpolation timestep, $q_{r}(A_{t}^{\tau}\mid A_{t})$ defines a reference trajectory (e.g., a Gaussian perturbation around a linear interpolation), and $u(\cdot)$ denotes the corresponding ground-truth velocity.
+Figure 3: Comparison of three paradigms: (a) Real-is-Sim, (b) Real-Only, and (c) Sim-to-Real, highlighting their online (real-world interaction) and offline (simulation-based) capabilities.
+Real-is-Sim offers a unified framework where deploying in a virtual environment is a simplified version of deploying in the real-world, lacking only real-time correction.
+This ensures seamless transferability.
+Real-Only is confined to real-world execution.
+Sim-to-Real struggles with distribution mismatch and dynamics discrepancies, making successful transfer uncertain.
+3 METHOD
+Real-is-sim is a framework that uses a simulator as a mediator between the real world and a learnt policy.
+For this approach to work, the simulator must continuously synchronize with the physical world by updating its state based on real-world observations.
+In our work, we use Embodied Gaussians as the simulator, as it provides the necessary mechanisms for real-time alignment.
+This section first formalizes the framework, then details the complete pipeline: creating a scene, collecting demonstrations, training and deploying a policy, and evaluating the resulting policy.
+3.1 Framework
+Our framework consists of three subsystems: (i) the real-world system, (ii) the simulator, and (iii) the learned policy.
+Real-is-sim distinguishes between two modes: online, where the simulator is connected to and synchronized with the real-world system, and offline, where the system runs entirely in simulation and can be parallelized.
+We illustrate the connections in Figure 3a and  Figure 4.
+The real-world system is composed of a set of cameras that output RGB images, denoted as $o_{\text{rgb}}$, and a high-impedance joint controller that receives as input a desired joint configuration $c\in\mathbb{R}^{d}$ for a robot with $d$ degrees of freedom.
+The simulator maintains an environment state $s_{t}$, a simulated robot state $r_{t}$, an environment corrective input $u_{t}$, and a simulated robot control input $q^{\prime}_{t}$ at each timestep $t$.
+These are defined as
+$s_{t}=\{p^{i}_{t},R^{i}_{t},v^{i}_{t},\omega^{i}_{t}\}_{i=1}^{N_{\text{obj}}},\quad u_{t}=\{f^{i}_{t},\tau^{i}_{t}\}_{i=1}^{N_{\text{obj}}},\quad r_{t}=\{q_{t},\dot{q}_{t}\}.$
+Figure 4: Information flow between subsystems.
+For each simulated object $i\in{1,\dots,N_{\text{obj}}}$, the state is described by the position $p\in\mathbb{R}^{3}$, orientation $R\in SO(3)$, linear velocity $v\in\mathbb{R}^{3}$, and angular velocity $\omega\in\mathbb{R}^{3}$.
+The corrective control inputs, which align the simulated state with the real world, are computed as virtual forces $f\in\mathbb{R}^{3}$ and torques $\tau\in\mathbb{R}^{3}$ applied to each object.
+For the simulated robot, the state includes the joint positions $q\in\mathbb{R}^{d}$, joint velocities $\dot{q}\in\mathbb{R}^{d}$, and the desired joint configuration $q^{\prime}\in\mathbb{R}^{d}$.
+The next states $s_{t+1}$ and $r_{t+1}$ are computed through the simulator’s physics step: $s_{t+1},r_{t+1}=F(s_{t},u_{t},r_{t})$.
+If the real-world system is connected, the robot’s desired joint configuration $c$ is set to $q_{t}$ at each timestep and the corrective inputs $u_{t}$ are set to $u_{t}=C(s_{t},o_{\text{rgb}})$, where $C$ represents the corrective mechanism in Embodied Gaussians (illustrated in Fig.
+2).
+The policy is a function $\pi(r_{t},q_{t})$ that maps a representation $r_{t}=g(s_{t})$ to a trajectory $a_{t:t+H}\in\mathbb{R}^{m\times(d+1)}$, where $r_{t}$ is derived from the simulator state $s_{t}$, $m$ is the number of waypoints in the trajectory (also called the action horizon).
+Each waypoint $a$ consists of a desired joint configuration $q^{\prime}\in\mathbb{R}^{d}$ and the estimated progress towards task completion $p\in[0,1]$.
+Once the trajectory is computed, the simulated robot follows it over a time horizon $H$, updating its desired joint configuration $q^{\prime}_{\text{desired}}$ every $H/m$ seconds to reach the next waypoint in the trajectory $a$.
+After the trajectory is completed, a new one is generated.
+In our implementation, we set $H=1~{}\text{second}$ and $m=32$.
+3.2 Pipeline
+Scene Setup We construct the environment following the Embodied Gaussians procedure.
+Each object is generated from 4 viewpoints and is ultimately represented by a position $p$, a rotation $R$, and a set of fixed-radius spheres for collision geometry.
+Visual appearance is modeled by attaching Gaussians to each body, optimized using Gaussian splatting [2].
+The ground and robot geometry and appearance are manually specified.
+Demonstration Collection Demonstrations are collected with Embodied Gaussians continuously correcting the simulated state using live camera observations.
+Collecting with the simulator in-the-loop allows users to monitor synchronization quality and adapt their demonstrations to maintain it.
+To improve synchronization, during demonstrations we minimize occlusions by keeping the robot body out of camera views, move quickly when not contacting objects, slow down during interactions, and maintain ample clearance around object edges.
+Throughout each demonstration, we record the full sequence of simulation states at 60Hz.
+Although the simulation models rich physical properties, only the object poses and robot joint configurations change over time; all other physical parameters are initialized once and remain constant.
+Similarly, visual states (the Gaussians) are only stored at the initial timestep.
+A full list of static parameters is provided in [7].
+Figure 5: The two modes of the Real-is-Sim framework: real-world deployment and virtual environment deployment.
+Policy Training Our policy outputs an action trajectory $a_{t:t+H}$, where each waypoint consists of a desired joint configuration $q^{\prime}\in\mathbb{R}^{d}$ and a scalar progress value $p\in[0,1]$.
+The progress value represents the relative completion of the task and is labeled based on the elapsed time in the demonstration, assuming that all demonstrations end at the target position.
+This heuristic, validated in [8], allows us to evaluate policies without manually engineering rewards.
+The policy is conditioned on the robot’s current joint configuration $q_{t}$.
+For object state-based policies, we additionally condition on the object’s pose, represented by a position and quaternion [^1] The quaternion is used for simplicity and did not introduce ambiguity in our tasks..
+For image-based policies, we encode rendered images into 64-dimensional vectors using a ResNet.
+Embodied Gaussians enables the rendering of images from arbitrary viewpoints, allowing the creation of virtual cameras.
+The representations used for conditioning, including object pose and rendered images, are shown in Figure 6.
+Policy Deployment During deployment, the simulator continues running alongside the real robot to ensure it remains synchronized with the physical world.
+The trained policy processes the state of the simulator and outputs the desired joint configurations for the simulated robot.
+This means the policy is operating with the same type of inputs it was trained on, controlling the simulated robot as it was during demonstration collection.
+The integration with the real-world system occurs through two key mechanisms: (1) RGB images from the real-world cameras are continuously input into the system, providing real-time corrections to the state of the simulated objects.
+This feedback enables the simulator to continuously adjust and align the simulated environment with the real world.
+(2) The real robot’s joint controller is configured to follow the simulated robot’s joint positions $q$ (and not the desired joints $q^{\prime}$).
+This design choice, where the real robot acts as a follower, allows for a smooth transition between operating on a real robot and operating solely in a virtual environment.
+Policy Evaluation By deactivating the real-world coupling and corrective force calculation ($u_{t}$), the same setup can be used in a virtual-only environment.
+This allows for easy parallelization of environments which enables fast policy evaluation.
+In virtual deployments, we achieve a speedup of 5x to 8x, depending on the chosen representation for the policy, when running 20 environments concurrently.
+Figure 6: Comparison of representations from the Real-to-Sim framework versus a real image baseline: (a) State-based representation (object position and orientation), (b) Virtual camera on robot end effector, (c) Virtual camera at real camera position, (d) Real camera image, which is used to correct the simulation, but not as a direct input to the policy.
+4 EXPERIMENTS
+Figure 7: Success rate of an imitation learning policy versus training steps, comparing evaluation on the real setup and in a virtual environment.
+The consistency between the two curves shows that the faster virtual-only evaluation is a reliable proxy for real-world testing, providing a practical alternative for policy assessment.
+Error bars represent 95% confidence intervals.
+Experimental Setup We evaluate our framework on the PushT task: a long horizon task where a robot uses a pusher to position a T-shaped object into a predefined location.
 The T-Block’s geometry (represented as a set of spherical shapes attached to a single body frame) is automatically generated using the method from Embodied Gaussians [1].
@@
 Gaussians attached to the robot, table, and T-Block are learned once using the Gaussian Splatting procedure outline in [1] and reused across experiments.
-For state-based policies, we use the pose of the object encoded as a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\in\mathbb{SO}(3)$, where $\mathbf{t}=(x,y,z)$ and $\mathbf{q}=(q_{x},q_{y},q_{z},q_{w})$.
-For image-based policies, we use a 64 dimensional feature vector output by a Resnet [22].
+For object state-based policies, we use the pose of the object encoded as a translation vector $\mathbf{p}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{R}\in\mathbb{SO}(3)$.
+For image-based policies, we use a 64 dimensional feature vector output by a ResNet [5] which was jointly trained with the policy.
 The images input to the ResNet are rendered from the simulation using Gaussian Splatting from virtual cameras.
@@
 In our experiments, we use a static camera placed in the same location as the real camera and a virtual gripper camera which we mount to the simulated robot’s end-effector (these are shown in Figure 6).
-To measure policy success on the PushT task, we define 20 in-distribution starting poses on the table that were not part of the demonstration set.
-In real-world testing, success is determined by whether the policy drives the T-Block to the target location from these poses.
+To evaluate policy success on the PushT task, we define 20 starting poses on the table that are not part of the demonstration set.
+In real-world testing, success is determined by whether the policy moves the T-Block to the target location from these poses.
 Each pose is tested three times, resulting in 60 measurements per success rate evaluation.
-Error bars in figures represent 95% confidence intervals.
-V-B Offline Evaluation
-In this experiment, we demonstrate how real-is-sim’s offline mode facilitates policy training and evaluation.
+In virtual-only testing, success is determined when the policy outputs a progress value above 0.9, a strategy verified in [8], which serves as a good proxy for success without requiring engineered reward values.
+Offline Evaluation In this experiment, we demonstrate how real-is-sim’s offline mode facilitates policy training and evaluation.
 A key challenge in imitation learning lies in selecting optimal checkpoints for deployment.
@@
 For a state-based policy trained on 30 demonstrations, we implement a flow-matching policy trained for 50,000 steps (Figure 7).
-We then compare two evaluation methodologies: (i) offline rollouts using real-is-sim and (ii) real-world online evaluations.
-Our offline approach initializes environments with identical test cases to online evaluations.
-The offline evaluation calculates success rates through the policy-reported progress.
+We then compare two evaluation methodologies: (i) evaluation on only the virtual world (without the real robot attached and the corrective mechanisms engaged) and (ii) real-world online evaluations.
+Our virtual-only approach initializes environments with identical test cases to the real-world evaluations.
+The virtual-only evaluation calculates success rates through the policy-reported progress.
 While we use this metric to maintain generality, practitioners could implement more sophisticated reward functions for their specific tasks.
-We selected four representative checkpoints (10k, 15k, 25k, 50k) for online evaluation (each requires around 1 hour to perform) while evaluating more comprehensively through offline mode.
-Figure 5 reveals strong correlation between evaluation curves, demonstrating that faster offline evaluations (minutes vs.
-hours) effectively approximate real-world results.
+We selected four representative checkpoints (10k, 15k, 25k, 50k) for real-world evaluation and performed more comprehensive analysis using faster offline evaluation.
+Figure 7 shows that, while absolute performance may differ, the relative ordering of checkpoints is preserved - making offline evaluation a practical tool for selecting high-performing policies in minutes rather than hours.
 Notably, later checkpoints do not consistently outperform earlier ones, suggesting practitioners could use offline evaluation for either early stopping or optimal checkpoint selection.
 While this experiment uses identical initial states for both evaluation modes to establish correlation, real applications could benefit from randomized initial states during training-phase evaluations.
-V-C Offline Data Collection and Augmentation
-In this experiment, we demonstrate how real-is-sim’s offline mode can be used to to collect additional demonstrations to increase task performance (Figure 6).
+Figure 8: (a) Success rates of different demonstration collection strategies.
+(b) Success rates of policies using different representations to condition the flow matching policy.
+Error bars are 95% confidence intervals.
+Offline Data Collection and Augmentation In this experiment, we demonstrate how real-is-sim’s offline mode can be used to to collect additional demonstrations to increase task performance (Figure 8).
 We train a state-based policy using 30 real-world demonstrations collected in online mode - which achieves a success rate of around 57%.
@@
 Interestingly, this approach yields a similar performance improvement as the augmented dataset, underscoring the effectiveness of real-is-sim’s offline mode in efficiently enhancing policy training.
-V-D Representation Flexibility
-Another major benefit of real-is-sim’s offline mode is its flexibility in extracting diverse representations from the Embodied Gaussians simulation.
+Representation Flexibility Another major benefit of real-is-sim’s offline mode is its flexibility in extracting diverse representations from the Embodied Gaussians simulation.
 For instance, it enables access to privileged information—such as precise object poses—that would be difficult to obtain in the real world.
 Additionally, virtual cameras can be freely positioned within the system, facilitating visuomotor policy learning from multiple perspectives.
-In our experiments, we demonstrate that policies leveraging different representations—including state-based object poses, a gripper-mounted virtual camera, and a static virtual camera—all successfully solve the task (Figure 7, 4).
+In our experiments, we demonstrate that policies leveraging different representations—including state-based object poses, a gripper-mounted virtual camera, and a static virtual camera—all successfully solve the task (Fig.
+8b, Fig.
+6).
 Notably, these approaches achieve performance comparable to, and in some cases surpassing, the baseline policy that relies solely on real-world images and robot actions.
@@
 This flexibility allows us to identify the optimal representation for a given manipulation task in offline mode and seamlessly deploy it in online execution.
-VI Limitations
-While our framework demonstrates promising results, its applicability is inherently constrained by the capabilities of the underlying physical simulator.
-Scenarios that deviate significantly from the simulator’s modeling assumptions pose challenges to both model accuracy and corrective mechanisms.
-To address these constraints, future improvements could focus on enhancing simulator fidelity through adaptive parameter estimation (e.g., real-time optimization of physical properties like friction or mass, as explored in works such as [20] and  [30]).
-While this work focused on the PushT task to validate core contributions, the methodology is extensible to broader manipulation domains.
-Enabling applications such as multi-object rearrangement, cloth manipulation, or non-prehensile pushing will require tailoring the simulator’s dynamics model and corrective policies to handle domain-specific interactions.
-VII Future Work
+5 RELATED WORK
+Limitations of Real-World Policy Evaluation: Evaluating robotic policies directly in the real world, while reliable, suffers from being labor-intensive, time-consuming [11, 12], and facing scalability bottlenecks due to manual processes [13].
+Consequently, researchers explore simulation for safer, repeatable, and scalable evaluation [14, 15, 16, 17], often focusing on improving simulation fidelity [14, 17, 16, 15].
+Our framework differs from this approach by integrating a continuously synchronized simulator throughout the entire policy lifecycle (training, evaluation, deployment).
+This real-time alignment via sensor data enables direct deployment of simulation-trained policies without further tuning and ensures a unified observation and control space for scalable evaluation.
+Addressing the Sim-to-Real Gap in Robotic Manipulation: Simulation offers a scalable alternative to real-world evaluation [11], but the sim-to-real gap remains a key challenge.
+Existing strategies include domain randomization to enhance robustness [18] and domain adaptation to align representations [19], as well as methods focusing on aligning simulation dynamics [20, 21] or using real-to-sim-to-real approaches [22, 23].
+Frameworks like DeepMind’s robot soccer also combine neural rendering with domain randomization [24].
+While these methods aim to bridge the sim-to-real gap by making policies more robust or adapting representations, our ”real-is-sim” paradigm eliminates this gap entirely.
+By continuously aligning the simulator with the real world and treating it as the execution environment, we decouple the policy from sim-real discrepancies.
+The challenge of domain alignment shifts to the simulator itself, which is addressed through real-time updates.
+Leveraging World Models for Robotics: Learned world models enable robots to predict future outcomes for planning and control.
+These range from 2D predictors [25] to more structured 3D physically grounded models, such as those utilizing Gaussian Splatting [26, 1], which allow for real-time tracking and simulated interaction.
+Our work builds upon this by exploring the integration of such advanced 3D world models within a behavior cloning framework.
+This aims to harness the advantages of simulation directly in real-world tasks by training policies on the dynamically updated, real-time digital twin provided by these models.
+6 CONCLUSIONS
+We introduced real-is-sim, a unified framework for policy training, evaluation, and deployment that maintains a continuously corrected simulation throughout the entire pipeline.
+By treating simulation as a persistent, real-time interface, enabled by Embodied Gaussians, we eliminate the disconnect between training and deployment domains.
+This reframes policy evaluation as a simulation-native process, even in real-world settings.
+Our results on the PushT task show strong alignment between offline and real-world performance, demonstrating that real-is-sim enables scalable, efficient evaluation without sacrificing fidelity.
+We hope this framework serves as a foundation for future real-world learning algorithms that leverage this tight coupling between simulation and reality.
+7 Limitations
+While our framework demonstrates promising results, its applicability is fundamentally constrained by the fidelity of the underlying physical simulator.
+Scenarios that diverge significantly from the simulator’s modeling assumptions present challenges for both forward prediction and corrective mechanisms.
+Future work could address these limitations by improving simulator accuracy through adaptive parameter estimation—for example, real-time optimization of physical properties such as friction or mass, as explored in [26] and [27].
+Another limitation of the current framework is its reliance on visual forces to correct the simulator state.
+This restricts applicability to scenarios where objects are frequently visible and do not undergo prolonged occlusion or manipulation without observation.
+Moreover, for Embodied Gaussians to function effectively, visual corrections must not conflict significantly with the dynamics predicted by the physics simulator.
+The approach assumes that both the simulator and the visual feedback are aligned, that is, they attempt to move the object in compatible directions.
+When this assumption breaks down and the corrections diverge, synchronization between the simulation and the real world may not be reliable.
+However, this limitation can be addressed.
+The synchronization mechanism can be extended with additional corrective signals beyond visual forces.
+For example, semantic cues, such as correspondence detection via methods like TAPIR [28] or loss augmentation using segmentation masks as in [29], could provide more robust alignment in challenging settings.
+The current framework was designed to demonstrate the simplest viable instantiation.
+Future extensions could incorporate these more sophisticated correction signals to broaden applicability across diverse and visually ambiguous environments.
+Another limitation lies in the scene initialization process.
+In this work, we employed a simple scheme that is sensitive to factors such as the number and placement of cameras, object complexity, and visibility.
+However, recent advances in learned priors for geometry initialization, such as InstantMesh [30] and InstantSplat [31], offer promising alternatives that could replace or enhance our approach, expanding the set of objects that can be initialized effectively.
+Although we focused on the PushT task to validate the core contributions, the proposed methodology is broadly applicable to other manipulation domains.
+Extending this approach to tasks such as multi-object rearrangement, cloth manipulation, or non-prehensile pushing will require adapting the simulator’s dynamics and corrective policies to accommodate more complex interactions.
+ACKNOWLEDGMENT
+The QUT authors acknowledge continued support from the Queensland University of Technology (QUT) through the Centre for Robotics.
+Their work was partially supported by the Australian Government through the Australian Research Council’s Discovery Projects funding scheme (Project DP220102398).
+Supplementary Material
+Real-is-Sim: Bridging the Sim-to-Real Gap with a Dynamic Digital Twin
+Appendix A Follower Configuration vs Leader Configuration
+We present real-is-sim as a framework that employs a continuously corrected digital twin of the environment to decouple a behavior cloning policy from the real robot.
+A central design decision in this framework is to configure the real robot as a follower of the simulated robot.
+In this setup, when the real robot is connected, it directly mirrors the joint states of the simulated counterpart.
+This follower design introduces a distinctive characteristic: the real world consistently lags slightly behind the simulated twin.
+To our knowledge, this temporal asymmetry is unique to Real-is-Sim and implies that the system always possesses a short-horizon preview of future states.
+While we find this property conceptually interesting, our primary motivation for adopting the follower configuration is practical, it results in a significantly simpler software architecture and reduces the complexity of sim-to-real integration.
+Supplementary Figure 1 illustrates this distinction.
+In both offline and online modes under the follower configuration (our proposed setup), the system interconnections (blue lines) remain constant.
+In contrast, the leader configuration—where the real robot dictates state and the simulator reacts—requires dynamic changes to the system architecture.
+Specifically, three data streams must cross the sim-to-real boundary in the leader case:
+  1. 1.
+    Control inputs from the policy to the real robot,
+  2. 2.
+    Control inputs from the real robot to the simulator,
+  3. 3.
+    Observations from the real robot to the simulator.
+By comparison, the follower configuration requires only two such streams.
+This reduction in cross-boundary communication simplifies the overall system and motivates our recommendation of the follower configuration in Real-is-Sim.
+Appendix B Benefits of Collecting Demonstrations with the Real Robot
+Figure 2: Tracking failure in two policies, one trained on demonstrations collected with the real robot in the loop, and the other trained entirely in simulation.
+The left image illustrates a failure case where the simulated physics conflicts with the visual correction required to stay in sync with the real world.
+Since physics acts as a hard constraint, the correction cannot be applied.
+Collecting demonstrations in online mode helps avoid such failure modes by encouraging behaviors that maintain alignment between the simulation and reality.
+In real-is-sim, we advocate collecting demonstrations with the real robot attached (online mode).
+During data collection, the user observes the state of the simulator overlaid with the real world, allowing them to continuously assess synchronization quality.
+This enables the user to adapt their behavior not only to complete the task but also to maintain alignment between the simulation and the real environment.
+Practically, this means knowing when to move quickly or slowly, and how much clearance to give during fast motions to accommodate slight mismatches between the two domains.
+Supplementary Figure 2 illustrates an exaggerated case in which demonstrations are collected entirely in simulation and the resulting policy is deployed on the real robot.
+Although the task is successfully demonstrated in simulation, it is done with insufficient clearance around objects—causing impacts in the virtual world that do not occur in the real one.
+The perception system’s correction mechanisms are unable to recover from this divergence, leading to a marked increase in tracking errors during real-world execution.
+For this reason, we recommend that at least some demonstrations be collected with the real robot in the loop.
+This allows both the policy and the demonstrator to implicitly learn behaviors that help keep the simulator and real world in sync.
+Appendix C Implementation Details
+We use an NVIDIA RTX 6000 Ada as our primary compute device.
+The robot employed is a Franka Emika with 7 degrees of freedom.
+The high-impedance controller for the real robot utilizes Ruckig [32] to generate joint trajectories with minimal jerk, in conjunction with the Franka Emika’s built-in joint controller.
+During demonstration collection, the simulated robot is constrained to act within a 2D plane using Pink [33].
+Note that this 2D Cartesian controller is only active during demo collection and only controls the simulated robot.
+The real robot continues to execute a joint impedance controller based on the simulated robot’s joint positions.
+Interestingly, the policy learns to operate in 2D, even though the joint controller does not impose such a restriction when the policy is deployed.
+Appendix D Alterations to Embodied Gaussians
+Our adaptation of Embodied Gaussians [1] (built using [7]) differs from the original framework in two key aspects.
+First, we omit shape-matching constraints for deformable objects, modeling all entities as rigid bodies composed of spherical collision primitives.
+This results in a rigid body simulator rather than a particle-based one, improving performance from 30Hz to 60Hz in online mode.
+Second, we treat the robot as a dynamic entity, simulating mass and inertia for its links and joints, in contrast to the original kinematic formulation, where robot motion was purely kinematic.
+This change enhances simulation stability and allows the simulated robot to react to the environment.
+We also found it essential to ensure that the bodies comprising the simulated robot are not affected by gravity.
+Without this, controlling the simulated robot becomes more difficult, as the controller would need to include a gravity compensation term.
+Appendix E Future Work
 While the always-in-the-loop simulator already corrects itself using observations, its accuracy could be further enhanced by backpropagating into physics parameters or learning a residual physics network, reducing reliance on explicit corrections by making the simulator’s predictions more physically faithful.
@@
 This would enrich policy training and further make use of the simulator’s role as a mediator between the real world and the policy.
-VIII CONCLUSIONS
-We introduced real-is-sim, a unified framework for policy training, evaluation, and deployment that maintains a continuously corrected simulation throughout the entire pipeline.
-By treating simulation as a persistent, real-time interface, enabled by Embodied Gaussians, we eliminate the disconnect between training and deployment domains.
-This reframes policy evaluation as a simulation-native process, even in real-world settings.
-Our results on the PushT task show strong alignment between offline and real-world performance, demonstrating that real-is-sim enables scalable, efficient evaluation without sacrificing fidelity.
-We hope this framework serves as a foundation for future real-world learning algorithms that leverage this tight coupling between simulation and reality.